OpenStudy (anonymous):

3. Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (

OpenStudy (anonymous):

do u need help?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

did u try?

OpenStudy (anonymous):

yes but it is confuse and i do virtual school so its explanation on how to do it was not easy to get

OpenStudy (anonymous):

assume that the number of dimes = X and the number of quarters = Y

OpenStudy (anonymous):

okay i wrote it down

OpenStudy (anonymous):

so X+Y=46 from that we ca say that X=46-Y

OpenStudy (anonymous):

is that clear to you?

OpenStudy (anonymous):

yes kind of

OpenStudy (anonymous):

what's not clear?

OpenStudy (anonymous):

well im still not sure how to solve it

OpenStudy (anonymous):

we still need a second equation to solve the problem

OpenStudy (anonymous):

so would i put the equations together to make the second one?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

the second one is as follow: 10X+25Y=700

OpenStudy (anonymous):

DO U KNOW WHY?

OpenStudy (anonymous):

because the quarter is 25 and the dime is 10 and the total is 7.00

OpenStudy (anonymous):

the dime = 10 cent the quarter = 25 cent $7=700 cent

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so now put them together?

OpenStudy (anonymous):

substitute X=(46-Y) into the second equation to find Y

OpenStudy (anonymous):

10(46-y)+25y=7.00?

OpenStudy (anonymous):

700 NOT 7.00

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

CAN U FIND Y?

OpenStudy (anonymous):

- 460 from both sides? or y ?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

-460

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got 240

OpenStudy (anonymous):

WHAT DID YOU GET?

OpenStudy (anonymous):

SO 15Y=240 Y=???

OpenStudy (anonymous):

y= 16

OpenStudy (anonymous):

X=?? USE THE FIRST EQUATION

OpenStudy (anonymous):

x = 36

OpenStudy (anonymous):

X=46-Y

OpenStudy (anonymous):

x= 30

OpenStudy (anonymous):

SO NUMBER OF DIMES = 30 NUMBER OF QUARTER = 16

OpenStudy (anonymous):

thank you i understood much better now

OpenStudy (anonymous):

30 DIMES + 16 QUARTERS = 700 CENTS OR $7.

OpenStudy (anonymous):

you welcome